# A Position Paper on the Automatic Generation of Machine Learning Leaderboards

### v1

@@ 1 Introduction
However, leaderboards on these platforms are often incomplete or missing for certain tasks, and they typically rely on manual updates. To reduce this manual effort, recent work has focused on automatically extracting experimental outcomes (referred to here as “tuples”) from research papers to populate leaderboards. We refer to this body of work as Automatic Leaderboard Generation (ALG): “A systematic process for extracting relevant experimental findings from scientific publications to create and maintain a leaderboard.”. Figure 3 illustrates an example of this process, showing the extraction of task, dataset, metric, score tuples from two research papers to construct a leaderboard.
@@ 2 Overview of Problem Definition
The ALG field has seen many advances over the years. At a broad level, the ALG task is an information extraction task, to extract a tuple containing key details of an ML experimental result.444 We acknowledge that ALG work rests on a long history of work in information extraction (IE) in scientific literature. The full body of IE work is out of scope for this analysis but is introduced briefly in Appendix A.
@@
A key issue in the field is the variation in input and output expectations across studies. Table 1 lists key ALG papers we examined, focusing on recent work using transformer models that enable data scaling.555Details on prior work are in Appendix C.
We can characterise the key differences in the problem definition as concerning expectations about input and output data. Specifically, we discuss: (1) reliance on domain knowledge, and (2) limited scope of extraction.666We also note that various works have differed in expectations on the data format (e.g., PDF or LaTeX). However, we do not see this as critical in hindering comparisons of results.
@@ 2.1 Reliance on Domain Knowledge
We observe that the ALG domains can be categorised as having different levels of reliance on prior domain knowledge, which ultimately impacts what information can be extracted. Essentially, two variants of the problem have been previously tackled: closed domain and open domain.777The “open domain” category includes hybrid cases that start with no domain knowledge and incrementally builds up knowledge as publications are processed.
@@
The closed-domain ALG problem stipulates that all the entities or tuples are predefined.888As in, bound by the closed world assumption. In the field, there have been two subvariants that we name: (1) predefined typed entities (PTE) and (2) predefined typed tuples (PTT).999We borrow “predefined” from sahinuc2024efficient.
@@
PTT is a further restriction beyond PTE in that only prescribed combinations of these science concepts are considered for establishing new tuples. We define PTT as “A closed-domain problem for ALG, in which a system is only allowed to detect leaderboard entries composed of specific, predefined combinations of known scientific concepts rather than forming any new combination.” In PTT variants of ALG, only predetermined combinations (often observed combinations) are used for creating new tuples (e.g., as in Hou2019Leaderboards).
@@ 2.2 Scope of Extraction
Furthermore, most work focuses only on extracting the top results from each paper, restricting each paper to a single entry per leaderboard Hou2019Leaderboards; Kardas2020AXCELL; Hou2021TDMSci; Yang2022TELIN. If a publication presents two methods, only the top-performing one typically appears on the leaderboard. This can lead to an incomplete and potentially biased view, omitting valuable contributions such as negative results.101010E.g., one may wish to compare neural networks with other machine learning methods (e.g., logistic regression, random forests) to evaluate the cost-benefit trade-off.
@@ 3 Overview of ALG Datasets
| Dataset | First Reported In | Variant | T | D | M | S | Md | LaTeX | HA | PwC | NLPP | Ann. | |
@@
With the growth of the field, several datasets have been proposed to evaluate ALG methods, making it hard for researchers to identify which datasets are best suited for benchmarking. To guide dataset selection, Table 2 summarises their key characteristics111111A more detailed version of this table can be found in Appendix D Table 5. We highlight the main dimensions along which datasets differ. The main takeaway from this table is the diversity of the datasets that have been used in past research, making it hard to make fair comparisons. We discuss the variations below. A few recent datasets offer valuable attributes: LEGOBench Singh2024LEGOBENCH is the largest and covers the broadest tuple scope (including score), while SciLead sahinuc2024efficient stands out for its exhaustive manual annotations.
@@ 3.1 ML Experiment Science Entities
As prior work has varied in the entity classes studied, datasets have likewise differed in the scope of their tuple and entity annotations. The most common format is task, dataset, metric, score (NLP-TDMS, Hou2019Leaderboards, PwC-LB Kardas2020AXCELL, TDMS-Ctx kabongo2024effective, SciLead sahinuc2024efficient), while the most comprehensive format is task, dataset, metric, score, method (LEGOBench, Singh2024LEGOBENCH). These five datasets can be considered “complete” leaderboard datasets, as they include the score within the tuple.131313These datasets can sometimes be divided into further subsets based on the size of the leaderboard. E.g., the ORKG-PwC and NLP-TDMS datasets filter out leaderboards with less than five entries. Datasets can also be divided into pre-defined subsets. E.g., the ORKG datasets include pre-defined splits that correspond experimentation by kabongo2024effective121212zero-shot versus few-shot, and the type of document representation: DocREC, DocTAET and DocFull. In contrast, two related datasets do not include scores (ORKG-PwC Kabongo2021Mining, SciREX Jain2020SCIREX.141414Although the paper does not mention recording the score, we found that the Github dataset includes a score. It is unclear whether this was added after the publication of the paper. https://github.com/allenai/SciREX
@@ 3.2 Source of Annotations
Most datasets are assembled using manually curated leaderboards as a distant supervision source. For example, the first leaderboard dataset, NLP-TDMS Hou2019Leaderboards, was derived from a community-maintained GitHub repository NLP Progress151515https://github.com/sebastianruder/NLP-progress, tracking state-of-the-art NLP datasets and tasks. With the growing popularity of Paper with Code, many researchers turn to this resource to build ALG datasets, including ORKG-PwC, PwC-LB, SciREX, TDMS-Ctx and LEGOBench.
@@ 3.3 Format of the Papers
Datasets differ in publication formats. PDFs, though common, mix presentation with logical structure, whereas LaTeXoffers cleaner organisation. Some datasets use only one format, PDF (LEGOBench, SciLead) or LaTeX(TDMS-Ctx), while others provide both (NLP-TDMS, ORKG-PwC, PwC-LB). We note that this distinction is less important as tools like Grobid lopez2009grobid grow in maturity to transform PDF files into a logical structure format, such as XML.
@@ 4 Overview of ALG Evaluation Metrics
One key issue in the field has been the use of various metrics for ALG evaluation, hindering result comparisons. Appendix E lists all metrics used in leaderboard experiments. Below, we outline the key evaluation metrics used in prior work.
@@ 4.1 Precision, Recall and F1
Although not explicitly stated, we believe that generally these scores are calculated per paper and then averaged. However, Singh2024LEGOBENCH calculated precision and recall per leaderboard. Experimental results can vary significantly depending on whether metrics are averaged across papers, leaderboards, or entities/tuples. To demonstrate this significance, we replicated an experiment of sahinuc2024efficient and found that if authors had used global averaging instead of per paper averages the recall would differ by 12.61.161616The authors conducted a zero-shot experiment evaluated using exact match. They reported a recall of 47.53 when averaging per paper, whereas the recall would have been 34.92 if averaged globally across all tuples. In Table 6 (Appendix E), we provide definitions of these metrics.
@@ 4.3 Granularity of Science Concepts
As science advances, scientific concepts evolve. For example, broad terms like neural LMs may split into finer categories (e.g., pre-trained LMs vs. LLMs), or sibling concepts may merge or become unevenly prominent (e.g., abstractive summarisation overtaking extractive summarisation with generative AI). Relatedly, capturing fine-grained method attributes, such as hyperparameters for neural networks, becomes increasingly important.
@@ 5.1 Document Representation
We note that several papers focus on finding the best representation of paper contents, whether starting from PDF or structured formats like LaTeXõr XML. Such representations help highlight key information, especially when later ML components must process limited input text.
@@ 5.2.2 Regarding Open Domain Approaches
In sahinuc2024efficient, detected entities correspond to concepts that fall into two categories: (1) unseen (i.e., new) and (2) seen. Using a leaderboard database that is initially empty, entities are checked for corresponding entries, with either an exact match or a partial match. If a match exists, the existing form in the database is used as the canonical representation for that concept. This can be viewed as a data normalisation step. For all unmatched entities, these are treated as unseen, and a new database entry is created for it.
@@ 5.2.3 A Note on Score Extraction
Despite being central to ALG, only a handful of works Hou2019Leaderboards; Kardas2020AXCELL; Singh2024LEGOBENCH; kabongo2024effective; sahinuc2024efficient extract best scores. Other work focused on extracting the experimental conditions. We note that this is a precursor to finding the full tuple for ALG (identifying experimental conditions to which the best score belongs). For works that extract best scores, methods vary. Hou2019Leaderboards apply heuristics based on orthographic features (boldface), whereas Kardas2020AXCELL use more complex inferences, classifying table cells as numeric or non-numeric. Extracted quantities are normalised and the extreme (maximum or minimum) score is kept based on the metric. Earlier models used dedicated methods to align scores with conditions, whereas recent LLM prompting extracts entire tuples, including scores, with a single task-based prompt kabongo2024effective; Singh2024LEGOBENCH; sahinuc2024efficient.
@@ 5.3 Tuple Verification and Entity Alignment
For each extracted tuple, the system should verify its correctness, especially for LLM-based approaches, which risk hallucinations. Pre-LLM methods often implicitly included this step within the extraction process. For example, by framing the tuple generation task as an NLI problem, Hou2019Leaderboards extract tuples that are aligned with the source content and entailed by the source text, essentially performing verification. Others use partial alignment of the tuple at the entity level, such as using a Bayesian model to map different equivalent referring expressions to a canonical value Kardas2020AXCELL.
@@ 5.4 Updating Leaderboard Database
Once a tuple is verified, the final step is updating the leaderboard database. Kardas2020AXCELL link experimental conditions to existing Papers with Code entries. Data may be normalised prior to this step sahinuc2024efficient, and filtered to exclude, for example, ablation studies Kardas2020AXCELL. Most prior work does not detail this step, as the focus lies on NLP techniques for extraction rather than their downstream application, despite often being motivated by it.
@@ 6.2 Dataset Reporting
We recommend that researchers report results on publicly available datasets as a minimum requirement. We highlight SciLead and LEGOBench as two suitable options. SciLead is valuable for its fully human-curated annotations, ensuring high quality. LEGOBench offers the largest dataset with broad tuple coverage, enabling large-scale benchmarking across diverse tasks and methods. These two datasets are complementary: SciLead provides a gold standard for high-accuracy evaluation, while LEGOBench allows robust assessment at scale. The feasibility of achieving broader and more informative evaluations strongly depends on ensuring open access to such datasets. Fortunately, SciLead and LEGOBench are fully open-source and thus support the practical feasibility of standardised evaluation without subscription or copyright barriers. However, a limitation of both datasets is that they only cover a restricted set of metadata attributes and focus solely on extracting the best results per paper. Therefore, in the next section (§7.6), we recommend that researchers develop more comprehensive datasets that include all reported results and richer metadata.
@@ 6.3 Metrics
In open-domain settings, exact string matching may be overly restrictive. We recommend reporting partial match metrics, which account for fuzzy or approximate matches. Such metrics better capture performance when multiple valid surface forms exist for the same scientific concept. This reflects real-world feasibility more accurately.
@@ 7.1 New or Unseen Entities
The 2024 shared task on ALG d2024overview highlights that many aspects of the task are still unsolved. It includes closed and open domain subtasks, with the latter involving new entity detection.171717The organisers refer to these as few-shot and zero-shot, referring on current ML terminology. Indeed, Kabongo2023ZeroShot showed that ML performance in extracting tuples with new entities (i.e., new scientific concepts, such as a newly introduced ML task or dataset) is much lower than extracting tuples with previously observed entities. In production, a challenge will be the feasibility of canonicalisation and disambiguation of these newly introduced ML entities. New entities often have ambiguous and inconsistent naming. For example, a newly introduced dataset might be referred to in short and long forms or with typos. In practice, feasibility depends on having automated canonicalisation methods that can cluster or align different surface forms of unseen entities. Without this, leaderboard entries will fragment into inconsistent records, undermining usability.

### v2

@@ 1 Introduction
However, leaderboards on these platforms are often incomplete or missing for certain tasks, and they typically rely on manual updates. To reduce this manual effort, recent work has focused on automatically extracting experimental outcomes (referred to here as “tuples”) from research papers to populate leaderboards. We refer to this body of work as Automatic Leaderboard Generation (ALG): “A systematic process for extracting relevant experimental findings from scientific publications to create and maintain a leaderboard.”444All acronyms used in this paper are listed in Appendix A Table 3. Figure 3 illustrates an example of this process, showing the extraction of task, dataset, metric, method, score tuples from two research papers to construct a leaderboard.
@@ 2 Overview of Problem Definition
The ALG field has seen many advances over the years. At a broad level, the ALG task is an information extraction task, to extract a tuple containing key details of an ML experimental result.555We acknowledge that ALG work rests on a long history of work in information extraction (IE) in scientific literature. The full body of IE work is out of scope for this analysis but is introduced briefly in Appendix B.
@@
A key issue in the field is the variation in input and output expectations across studies. Table 1 lists key ALG papers we examined, focusing on recent work using transformer models that enable data scaling.666Details on prior work are in Appendix D.
We can characterise the key differences in the problem definition as concerning expectations about input and output data. Specifically, we discuss: (1) reliance on domain knowledge, and (2) limited scope of extraction.777We also note that various works have differed in expectations on the data format (e.g., PDF or LaTeX). However, we do not see this as critical in hindering comparisons of results.
@@ 2.1 Reliance on Domain Knowledge
We observe that the ALG domains can be categorised as having different levels of reliance on prior domain knowledge, which ultimately impacts what information can be extracted. Essentially, two variants of the problem have been previously tackled: closed domain and open domain.888The “open domain” category includes hybrid cases that start with no domain knowledge and incrementally builds up knowledge as publications are processed.
@@
The closed-domain ALG problem stipulates that all the entities or tuples are predefined.999As in, bound by the closed world assumption. In the field, there have been two subvariants that we name: (1) predefined typed entities (PTE) and (2) predefined typed tuples (PTT).101010We borrow “predefined” from sahinuc2024efficient.
@@
PTT is a further restriction beyond PTE in that only prescribed combinations of these science concepts are considered for establishing new tuples. We define PTT as “a closed-domain problem for ALG, in which a system is only allowed to detect leaderboard entries composed of specific, predefined combinations of known scientific concepts rather than forming any new combination.”
In PTT variants of ALG, only predetermined combinations (often observed combinations) are used for creating new tuples (e.g., as in Hou2019Leaderboards).
@@ 2.2 Scope of Extraction
Furthermore, most work focuses only on extracting the top results from each paper, restricting each paper to a single entry per leaderboard Hou2019Leaderboards; Kardas2020AXCELL; Hou2021TDMSci; Yang2022TELIN. If a publication presents two methods, only the top-performing one typically appears on the leaderboard. This can lead to an incomplete and potentially biased view, omitting valuable contributions such as negative results.111111E.g., one may wish to compare neural networks with other machine learning methods (e.g., logistic regression, random forests) to evaluate the cost-benefit trade-off.
@@ 3 Overview of ALG Datasets
| Dataset | First Reported In | Variants | T | D | M | S | Md | LaTeX | HA | PwC | NLPP | Ann. | |
@@
With the growth of the field, several datasets have been proposed to evaluate ALG methods, making it hard for researchers to identify which datasets are best suited for benchmarking. To guide dataset selection, Table 2 summarises their key characteristics.121212A more detailed version of this table can be found in Appendix E Table 6 We highlight the main dimensions along which datasets differ. The main takeaway from this table is the diversity of the datasets that have been used in past research, making it hard to make fair comparisons. We discuss the variations below. A few recent datasets offer valuable attributes: LEGOBench Singh2024LEGOBENCH is the largest and covers the broadest tuple scope (including score), while SciLead sahinuc2024efficient stands out for its exhaustive manual annotations.
@@ 3.1 ML Experiment Science Entities
As prior work has varied in the entity classes studied, datasets have likewise differed in the scope of their tuple and entity annotations. The most common format is task, dataset, metric, score (NLP-TDMS, Hou2019Leaderboards, PwC-LB Kardas2020AXCELL, TDMS-Ctx kabongo2024effective, SciLead sahinuc2024efficient), while the most comprehensive format is task, dataset, metric, score, method (LEGOBench, Singh2024LEGOBENCH). These five datasets can be considered “complete” leaderboard datasets, as they include the score within the tuple.141414These datasets can sometimes be divided into further subsets based on the size of the leaderboard. E.g., the ORKG-PwC and NLP-TDMS datasets filter out leaderboards with less than five entries. Datasets can also be divided into pre-defined subsets. E.g., the ORKG datasets include pre-defined splits that correspond to experimentation by kabongo2024effective131313zero-shot versus few-shot, and the type of document representation: DocREC, DocTAET and DocFull. In contrast, two related datasets do not include scores: ORKG-PwC Kabongo2021Mining, and, SciREX Jain2020SCIREX. Note that, for SciREX, the GitHub dataset includes a score.151515https://github.com/allenai/SciREX It is unclear whether this was added after the publication of the paper, demonstrating that data versioning can be a challenge.
@@ 3.2 Source of Annotations
Most datasets are assembled using manually curated leaderboards as a distant supervision source. For example, the first leaderboard dataset, NLP-TDMS Hou2019Leaderboards, was derived from a community-maintained GitHub repository NLP Progress161616https://github.com/sebastianruder/NLP-progress, tracking state-of-the-art NLP datasets and tasks. With the growing popularity of Paper with Code, many researchers turn to this resource to build ALG datasets, including ORKG-PwC, PwC-LB, SciREX, TDMS-Ctx and LEGOBench.
@@ 3.3 Format of the Papers
Datasets differ in publication formats of the source publications. PDFs, though common, mix presentation with logical structure, whereas LaTeX representations allows one to precisely isolate content from presentation. Some datasets use only one format, PDF (LEGOBench, SciLead) or LaTeX(TDMS-Ctx), while others provide both (NLP-TDMS, ORKG-PwC, PwC-LB). We note that this distinction is less important as tools like Grobid lopez2009grobid grow in maturity to transform PDF files into a logical structure format, such as XML.
@@ 4 Overview of ALG Evaluation Metrics
One key issue in the field has been the use of various metrics for ALG evaluation, hindering result comparisons. Appendix F lists all metrics used in leaderboard experiments. Below, we outline the key evaluation metrics used in prior work.
@@ 4.1 Precision, Recall and F1
Although not explicitly stated, we believe that generally these scores are calculated per paper and then averaged. However, Singh2024LEGOBENCH calculated precision and recall per leaderboard. Experimental results can vary significantly depending on whether metrics are averaged across papers, leaderboards, or entities/tuples. To demonstrate this significance, we replicated an experiment of sahinuc2024efficient and found that if authors had used global averaging instead of per paper averages the recall would differ by 12.61.171717The authors conducted a zero-shot experiment evaluated using exact match. They reported a recall of 47.53 when averaging per paper, whereas the recall would have been 34.92 if averaged globally across all tuples. In Table 7 (Appendix F), we provide definitions of these metrics.
@@ 4.3 Granularity of Science Concepts
As science advances, scientific concepts evolve. For example, broad terms like neural LMs may split into finer categories (e.g., pre-trained LMs vs. LLMs), or sibling concepts may merge or have their relative importance change (e.g., abstractive summarisation overtaking extractive summarisation as the dominant summarisation approach with the advent of deep learning). Similarly, what counts as an appropriate level of detail may change, such as the hyperparameters for neural networks, which has become part of standard reporting practice.
@@ 5.1 Document Representation
We note that several papers focus on finding the best representation of paper contents, whether starting from PDF or from structured formats like LaTeXor XML. Such representations help highlight key information, especially when later ML components must process limited input text.
@@
An alternative document representation that should be explored in future research for ALG is the use of image-based inputs, where PDF pages are treated as images and processed by vision-based or multimodal LLMs. This preserves layout structure crucial for accurate table interpretation, especially in cases where leaderboard-relevant results are embedded in complex visual formats. Recent OCR-free models khalighinejad2025matvix, demonstrate that bypassing textual conversion of scientific text can reduce noise and improve tuple extraction accuracy in visually rich scientific documents.
To support large-scale ALG deployment, document representations like DocTAET can also be used to reduce LLM input length and inference cost by filtering to only the most relevant sections (e.g., title, abstract, tables). This makes processing more efficient when applied across tens of thousands of papers.
@@ 5.2.2 Regarding Open Domain Approaches
In sahinuc2024efficient, detected entities correspond to concepts that fall into two categories: (1) unseen (i.e., new) and (2) seen. Using a leaderboard database that is initially empty, entities are checked for corresponding entries, with either an exact match or a partial match. If a match exists, the existing form in the database is used as the canonical representation for that concept. This can be viewed as a data normalisation step. For all unmatched entities, these are treated as unseen, and a new database entry is created for them.
For the ALG data normalisation step, we recommend caching normalisation decisions: once the LLM maps entity or tuple A to B, the same rule can be reused for identical cases, avoiding repeated LLM calls and reducing computational cost.
@@ 5.2.3 A Note on Score Extraction
Despite being central to ALG, only a handful of works Hou2019Leaderboards; Kardas2020AXCELL; Singh2024LEGOBENCH; kabongo2024effective; sahinuc2024efficient extract best scores. Other work focused on extracting the experimental conditions. We note that finding these conditions is a precursor to finding the full tuple for ALG (identifying experimental conditions to which the best score belongs). For works that extract best scores, methods vary. Hou2019Leaderboards apply heuristics based on orthographic features (boldface), whereas Kardas2020AXCELL use more complex inferences, classifying table cells as numeric or non-numeric. Extracted quantities are normalised and the extreme (maximum or minimum) score is kept based on the metric. Earlier models used dedicated methods to align scores with conditions, whereas recent LLM prompting extracts entire tuples, including scores, with a single task-based prompt kabongo2024effective; Singh2024LEGOBENCH; sahinuc2024efficient.
@@ 5.3 Tuple Verification and Entity Alignment
For each extracted tuple, the system should verify its correctness, especially for LLM-based approaches, which are susceptible to hallucination risk. Prior to the introduction of LLM, methods often implicitly included this step within the extraction process. For example, by framing the tuple generation task as an NLI problem, Hou2019Leaderboards extract tuples that are aligned with the source content and entailed by the source text, essentially providing some form of rationale for generated results. Others use partial alignment of the tuple at the entity level, such as using a Bayesian model to map different equivalent referring expressions to a canonical value Kardas2020AXCELL.
@@ 5.4 Updating Leaderboard Database
Once a tuple is verified, the final step is updating the leaderboard database. Kardas2020AXCELL link experimental conditions to existing Papers with Code entries. Data may be normalised prior to this step sahinuc2024efficient, and filtered to exclude, for example, ablation studies Kardas2020AXCELL. Most prior work does not detail this step, as the focus lies on NLP techniques for extraction rather than their downstream application, despite often being motivated by it. Efficient ALG database updates at scale require batched writes and schema-aware indexing. To avoid redundant updates, tuples can be deduplicated with hash-based checks.
@@ 6.2 Dataset Reporting
We recommend that researchers report results on publicly available datasets as a minimum requirement. We highlight SciLead and LEGOBench as two suitable options. SciLead is valuable for its fully human-curated annotations, ensuring high quality. LEGOBench offers the largest dataset with broad tuple coverage, enabling large-scale benchmarking across diverse tasks and methods. These two datasets are complementary: SciLead provides a gold standard for high-accuracy evaluation, while LEGOBench allows robust assessment at scale. The feasibility of achieving broader and more informative evaluations strongly depends on ensuring open access to such datasets. Fortunately, SciLead and LEGOBench are fully open-source and thus support the practical feasibility of standardised evaluation without subscription or copyright barriers. However, a limitation of both datasets is that they only cover a restricted set of metadata attributes and focus solely on extracting the best results per paper. Therefore, in Section §7.6, we recommend that researchers develop more comprehensive datasets that include all reported results and richer metadata. We also want to highlight the need for dataset versioning. The documentation ambiguity for SciREX Jain2020SCIREX, as discussed in §3.1, where the original paper omits the score while the GitHub repository later includes one, illustrates the problems which can arise when data differs from the original publication.
@@ 6.3 Metrics
In open-domain settings, exact string matching may be overly restrictive. We recommend reporting partial match metrics, which account for fuzzy or approximate matches. Such metrics better capture performance when multiple valid surface forms exist for the same scientific concept. This reflects real-world application scenarios more accurately.
@@ 7.1 New or Unseen Entities
The 2024 shared task on ALG d2024overview highlights that many aspects of the task are still unsolved. It includes closed and open domain subtasks, with the latter involving new entity detection.181818The organisers refer to these as few-shot and zero-shot, referring on current ML terminology. Indeed, Kabongo2023ZeroShot showed that ML performance in extracting tuples with new entities (i.e., new scientific concepts, such as a newly introduced ML task or dataset) is much lower than extracting tuples with previously observed entities. In production, a challenge will be the feasibility of canonicalisation and disambiguation of these newly introduced ML entities. New entities often have ambiguous and inconsistent naming. For example, a newly introduced dataset might be referred to in short and long forms or with typos. In practice, feasibility depends on having automated canonicalisation methods that can cluster or align different surface forms of unseen entities. Without this, leaderboard entries will fragment into inconsistent records, undermining usability.